# VidMuse: A Simple Video-to-Music Generation Framework with Long-Short-Term Modeling

### v1

@@ 5.4 Ablation Studies
Figure 5: Qualitative Comparison results on sound spectrograms produced by different methods.
Table 4: Ablation studies on model size.
     Models Params               Metrics
[…]
  VidMuse-S $455$M $0.854$ $57.317$ $4.695$ $1.349$ $0.433$ $0.165$
VidMuse-M $1.9$B 0.843 41.354 2.413 1.487 0.840 0.193
VidMuse-L $3.3$B $0.873$ $53.955$ $4.041$ $1.330$ $0.453$ $0.170$
5.5 Qualitative Analysis
Our qualitative analysis, presented in Fig. 5, underscores specific limitations in the methods employed by CMT and M2UGen. For CMT, the method’s reliance on extracting only low-level visual cues for music generation, particularly in creating symbolic music, leads to pronounced discontinuities. As illustrated in Fig. 5, “Abrupt gaps” occur, especially in darker video segments where the model fails to predict symbolic music notes, resulting in silence.
M2UGen’s use of LLMs and text embeddings to translate visual content into music affects its precision in capturing detailed visual nuances, leading to a noticeable lack of diversity, evidenced by repetitive musical themes. Additionally, M2UGen’s limitation to 30-second music generation segments requires the division of longer videos into clips for separate processing. This approach results in abrupt transitions between segments after stitching, which compromises the musical continuity.
The last row of Fig. 5 showcases our Long-Short-Term (LST) approach, highlighting its ability to produce globally diverse music that captures the essence of the ground truth. This method ensures our music generation remains contextually rich and aligned with the video.
6 Conclusion
Limitations. Our work achieves a significant advancement in video-to-music generation, but it still has some limitations. First, the current implementation relies on a EnCodec model [20] trained on a dataset that restricts the system’s capacity to a 32kHz audio sampling rate. This codec exhibits a noticeable reconstruction loss in the audio signal, potentially lowering the quality of the generated music. Second, training large models requires substantial computational resources. Our future work aims to overcome these limitations by integrating advanced codec technologies to enhance audio reconstruction fidelity, and exploring a more efficient method design.
Appendix A Additional Experiments
Visual Encoder. We conduct an ablation experiment to study the impact of visual encoders. As shown in table A2, we observe that VidMuse is robust in processing visual information for music generation. To trade-off between computational costs and quality of generation, we select CLIP [61] as the default visual encoder if not specified.
Codebook Pattern. The exploration of codebook interleaving patterns has attracted attention from researchers across several domains [90, 77, 11, 85, 41]. In our ablation study focusing on the patterns, we find that while the Parallel and Vall-E [77] patterns align with the findings for text-to-music generation in MusicGen [11], the flattened codebook pattern does not consistently exceed the performance of the delay pattern in tasks of generating music from video. The delay pattern, notable for its relatively low computational cost, is therefore selected for our implementation. The results of this study are presented in table A3.
Table A1: Ablation studies on the ratio of finetuning data.
     Finetuning Data             Metrics
[…]
  0 $0.750$ 35.788 $18.847$ $1.088$ $0.653$ $0.194$
10k $0.792$ $35.824$ $2.383$ $1.135$ $0.653$ $0.206$
20k 0.738 $36.171$ 2.369 $1.175$ $0.710$ 0.207
40k $0.771$ $36.521$ $2.476$ $1.150$ 0.716 $0.195$
60k $0.760$ $36.444$ $20.519$ 1.213 $0.706$ $0.197$
Table A2: Ablation studies on video encoders.
     Encoders               Metrics
[…]
  ViViT [6] $0.822$ $37.167$ $1.962$ $1.433$ $0.870$ $0.194$
VideoMAE [73] $0.778$ $37.900$ $19.388$ $1.074$ $0.663$ $0.209$
CLIP [61] $0.843$ $41.354$ $2.413$ $1.487$ $0.840$ $0.193$
ViT [19] $0.876$ $50.427$ $4.823$ $1.081$ $0.503$ $0.195$
(a) The distribution of video genres in our dataset.
(b) Comparisons with different datasets.
(c) Duration distributions of pretraining and finetuning data.
(d) Duration distribution of our benchmark data.
Figure A1: The statistics of our dataset.

### v2

@@ A.1 Additional Experiments
A.2 Details of Dataset Construction
Coarse Filtering. We design a rule-based filtering strategy for initial data screening. First, we perform illegal video and audio filters, which filter out the video without an audio track or a video track. Next, we apply a duration filter to filter out videos based on their duration, excluding those that are either too long (over 480 seconds) or too short (under 30 seconds). Additionally, we implement a domain filter to examine metadata and exclude specific categories such as Interview, News, and Gaming, which often have background music that lacks semantic alignment with the visual content. We also filter out videos containing inappropriate content, such as violence or explicit material.
Fine-grained Filtering. To further ensure the quality of our data, we conducted additional audio and visual analyses. For the audio analysis, raw videos may contain audio segments without music, such as speech, silence, etc. To ensure the final dataset consists of high-quality video-music pairs, we retain only those videos with a larger portion of music content. We utilize the sound event detection model PANNs (Kong et al., 2020), which provides frame-level event labels across the entire video to identify music events. Based on the observation from a subset of videos, we define two thresholds, i.e., a confidence threshold and a duration threshold, for analyzing the music event. The confidence threshold is set at $0.5$, indicating an audio frame is considered a music event if the PANNs model predicts the probability of the “Music” label to be over $0.5$. The duration threshold of a music event requires that at least $50\%$ of the audio’s frames are classified as music events for the video to be considered valid.
For the visual analysis, some videos only consisting of static images will be removed. Specifically, we uniformly sample multiple temporal windows without overlap from the video. Within each window, we use Structural Similarity Index Measure (SSIM) (Wang et al., 2004) between the first frame and the last frame. By aggregating average SSIM values from all temporal windows, we remove the videos with average SSIM values lower than a threshold of 0.8, empirically.
Music Source Separation. Since the irrelevant human speech in videos poses a negative impact on music generation, we apply music source separation to process the videos. We employ Demucs (Défossez et al., 2019) as the music source separation model to filter out the speech signals.
Audio-Video Alignment Ranking. ImageBind-AV (Girdhar et al., 2023) scores usually reflect the semantic correlation between the vision and audio modality. To construct a high-quality subset with better alignment, we compute the ImageBind-AV scores for all the data and rank them accordingly.
A.3 Details of Evaluation Metrics
Frechet Audio Distance (FAD) [^3] https://github.com/haoheliu/audioldm_eval (Kilgour et al., 2018) (Kilgour et al., 2018) is a reference-free evaluation metric for assessing audio quality. Similar to Frechet Image Distance (FID)(Heusel et al., 2017), it compares the embedding statistics of the generated audio clip with ground truth audio clips. A shorter distance usually denotes better human-perceived acoustic-level audio quality. However, this metric cannot reflect semantic-level information in audio. We report the FAD based on the VGGish(Hershey et al., 2017) feature extractor.
Frechet Distance (FD)33footnotemark: 3 measures the similarity between generated samples and target samples in audio generation fields. It’s similar to FAD but uses a PANNs feature extractor instead. PANNs(Kong et al., 2020) have been pre-trained on AudioSet(Gemmeke et al., 2017), one of the largest audio understanding datasets, thus resulting in a more robust metric than FAD.
Kullback-Leibler Divergence (KL)33footnotemark: 3 reflects the acoustic similarity between the generated and reference samples to a certain extent. It is computed over PANNs’ multi-label class predictions.
Density and Coverage [^4] https://github.com/clovaai/generative-evaluation-prdc (Naeem et al., 2020) measures the fidelity and diversity aspects of the generated samples. Fidelity measures how closely the generated samples match the real ones, while diversity assesses whether the generated samples capture the full range of variation found in real samples. We use CLAP(Wu et al., 2023b) embeddings for manifold estimation.
Imagebind Score [^5] https://github.com/facebookresearch/ImageBind (Girdhar et al., 2023) assesses to what extent the generated music aligns with the videos. Despite the fact that Imagebind extends the CLIP to six modalities, we only use the branches of audio and vision. Since we use ImageBind to filter out video-audio pairs with a low matching score during dataset construction, the ImageBind score is naturally used in our evaluation. We acknowledge that ImageBind is not specifically trained on music data, which may limit its effectiveness in capturing the full complexity of video-music alignment. However, it remains the most suitable option available for this task at present.
A.4 Details of the Inference Process
When predicting music on videos of arbitrary length, maintaining music consistency and coherence is particularly important. However, it leads to a significant challenge on computational resources due to the quadratic dependency of transformers-based models on sequence length (Zaheer et al., 2020; Beltagy et al., 2020). To address this problem, we adopt a sliding window approach for inferring the whole video.
During inference, given an input video with a length of $L$, we define $L_{s}$ as the length of the sliding window and $O$ as the overlap between consecutive windows. With the window start position $t$ initially set to 00, the inference involves the following steps compactly while $t+L_{s}\leq L$: (1) using a visual encoder to extract feature representations X and capture long-term dependencies $\textbf{X}_{l}$; (2) collecting embeddings within the window $[t,t+L_{s}]$ to obtain $\textbf{X}_{s}$; (3) predicting the music tokens $\bar{\textbf{Y}}$ for the reduced window $[t,t+L_{s}-O]$ based on $\textbf{X}_{l}$ and $\textbf{X}_{s}$; (4) decoding $\bar{\textbf{Y}}$ to the predicted audio $\bar{\textbf{A}}$ using the audio decoder; (5) move the window forward by setting $t=t+L_{s}-O$, and repeating steps (2) to (5) until the end of the video.
Figure A1: User study process. Participants evaluate the videos based on four criteria: Audio Quality, Video-Music Alignment, Musicality, and Overall Assessment.
Figure A2: Qualitative Comparison results on sound spectrograms produced by different methods.